OpenStudy (anonymous):

evaluate expression for m=2 and n=3 -6m+n

OpenStudy (anonymous):

If m = 2, I can replace 2 for m -6(2) + n If n =3, I can replace 3 for n -6(2) + 3 =-12+3 =-9 Does that make sense?